a 5 kg vrick of 20cm*10cm*8cm dimensionless lying on the largest base.it is now made to stand with length vertical.if g=10m/s calculate the amount of work done


a)3j


b)7j


c)5j


d)9j

Grade:11

1 Answers

KARAN ACHARYA
12 years ago

initially center of mass was at height 4cm from ground.

final height of center of mass is 10 cm

change in height = 10 -4 = 6cm

work done = change in potential energy

               =mgh [h is the change in height]

               =5x10x(6/100) =3j

 

PLEASE           APPROVE            MY             ANSWER           IF             YOU              LIKED               IT